Louisiana State University Health Sciences Center Administration & Finance
 

Friday, August 3, 2007

Dental School
  • Ceiling Installation
    • Landis continues to work of the punchlist items
    • 1st floor air handlers were disconnected and moved to their new temporary locations. Re-installation will take place next week.
    • Rough-in for light, fixtures and outlets continues.
  • Sterilizer
    • Door frame and sheetrock installation is complete.
  • Carpet Installation
    • Carpet Installation continues in Library
  • Remediation
    • Awaiting viable sample results
MEB DUCT CLEANING
  • Zones 4C and 4D are complete.
  • Unit 4-B was vacuumed, cleaned, and coated. The ducts will be coated this evening.

1542 Tulane

  • Ordes pulled wires from the disconnect, mounted at the entrance to the lot at the Medical School Building, to the guard booth in the L&M lot.
  • Siemens continues to remove old. fire alarm wiring and devices.
Landscaping Project
  • Installed last of the garden edging for Lion's Eye Building on the S. Prieur St. side.